# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 1. Introduction
For a positive integer $m$ and an arbitrary graph ${\cal G}$ (directed or undirected), the $m$-dimensional Weisfeiler-Leman algorithm (the $m$-dim $\operatorname{WL}$ algorithm) canonically constructs a special coloring of the $m$-tuples consisting of vertices of ${\cal G}$. Two given graphs are declared isomorphic if the resulting colorings do not distinguish these graphs; in this case, the graphs are said to be $\operatorname{WL}_{m}$-equivalent. The algorithm works correctly if the $\operatorname{WL}_{m}$-equivalence of the input graphs implies that they are isomorphic. The $m$-dim $\operatorname{WL}$ algorithm is the naive canonical vertex classification in the sense of [2] if $m=1$, and is the classical Weisfeiler-Leman algorithm [30] if $m=2$. As one of the key ingredient, the $m$-dim $\operatorname{WL}$ algorithm was used in the Babai quasipolynomial algorithm testing isomorphism of graphs [3].
Following [15], the Weisfeiler-Leman dimension ($\operatorname{WL}$-dimension) $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal G})$ of a graph ${\cal G}$ is defined to be the smallest positive integer $m$ such that any graph which is $\operatorname{WL}_{m}$-equivalent to the graph ${\cal G}$ is isomorphic to ${\cal G}$. The isomorphism of any graph of order $n$ and $\operatorname{WL}$-dimension $m$ with any other graph can be tested in time $O(n^{m+1}\log n)$. Now the question we started with can be reformulated as follows (see [28]): whether there exists a positive integer $M$ such that $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal G})\leq M$ for every circulant graph ${\cal G}$.
It is known that the $\operatorname{WL}$-dimension of a general graph cannot be bounded from above by an absolute constant [4]. On the other hand, it was proved in [28] that if ${\cal G}$ is a circulant graph of prime power order, then $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal G})\leq 3$. The main result of the present paper (Theorem 1.1 below) shows that the $\operatorname{WL}$-dimension of a general circulant graph can be arbitrary large. To formulate this result, we refer to Section 10 for the exact definition of $\varepsilon$-expanders, and given a positive integer $a$, denote by ${\mathfrak{B}}_{a}$ the set of all bipartite cubic $3$-connected graphs of order $a$.
@@
Following [32], we define the $\operatorname{WL}$-dimension of a circulant ${\cal G}$ with respect to the class of all circulant graphs as the smallest positive integer $m$ such that any circulant graph which is $\operatorname{WL}_{m}$-equivalent to ${\cal G}$ is isomorphic to ${\cal G}$. It was proved in [32] that this “relative” $\operatorname{WL}$-dimension of any circulant of order $n$ is at most $O(\log n)$. The question of the tight asymptotic for the maximum relative $\operatorname{WL}$-dimension of circulants remains open. As Corollary 1.2 shows, it is between $O(\sqrt{\log n})$ and $O(\log n)$.
@@
We conclude the introduction with a brief discussion of the proof of Theorem 1.1. Our technique is based on theory of coherent configurations [5], see also Section 2. If ${\cal G}$ is a graph, then the output ${\cal X}=\operatorname{WL}({\cal G})$ of the $2$-dim $\operatorname{WL}$-algorithm applied to ${\cal G}$ can naturally be viewed as a coherent configuration. The concept of the $\operatorname{WL}$-dimension extends to coherent configurations such that
$\dim_{\scriptscriptstyle\operatorname{WL}}({\cal G})=\dim_{\scriptscriptstyle\operatorname{WL}}({\cal X}),$
@@
see Section 4. When the graph ${\cal G}$ is circulant, the coherent configuration ${\cal X}$ naturally corresponds to a Schur ring ${\cal A}$ over a cyclic group (a circulant S-ring), see Section 5. Now to find a circulant graph ${\cal G}$ with large $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal G})$, we construct in Section 8 a circulant S-ring ${\cal A}^{\star}$ such that the corresponding coherent configuration ${\cal X}^{\star}$ has two properties: (a) $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal X}^{\star})$ is large, and (b) ${\cal X}^{\star}=\operatorname{WL}({\cal G})$ for some circulant graph ${\cal G}$. The construction of ${\cal A}^{\star}$ (see below) enables us to verify property (b) more or less easily, see Subsection 10.2.
@@
The construction of the S-ring ${\cal A}^{*}$ satisfying condition (a) is more subtle. First, we use a theory of multipliers in circulant S-rings, that was developed in [13, 14], see also Section 6. This theory is applied to express the $\operatorname{WL}_{m}$-equivalence of coherent configurations of circulant graphs in terms of $k$-local systems of multipliers of the corresponding circulant S-rings for $m<\sqrt{k}$, introduced and studied in Section 7. Thus our problem reduces to finding a circulant S-ring having a $k$-local system of multipliers for enough large $k$. An explicit construction of such an S-ring is given in Section 9 and is based on two ingredients defined in Section 8.
@@
The first ingredient is a class of coherent configurations, introduced and studied in [8], see Subsection 8.1. For any cubic graph ${\cal G}$, one can define a coherent configuration ${\cal X}={\cal X}({\cal G})$ of this class so that: if ${\cal G}$ has $a$ vertices, then ${\cal X}$ has $4a$ points, and if ${\cal G}$ is a $3$-connected $\varepsilon$-expander, then
(2) $\dim_{\scriptscriptstyle\operatorname{WL}}({\cal X})\geq c^{\prime}\cdot a$
@@
The second ingredient is a class consisting of special (circulant) S-rings ${\cal A}$ over cyclic groups of orders $n$ satisfying the conditions of Theorem 1.1. The structure of ${\cal A}$ depends only on the number $n$, and is enough simple to describe all possible multipliers. In particular, if we fix a bipartite graph ${\cal G}$ from the previous paragraph, then isomorphisms of the coherent configuration ${\cal X}$ yield multipliers of ${\cal A}$. Now the required S-ring ${\cal A}^{\star}$ is the algebraic fusion of the S-ring ${\cal A}$ with respects to the multipliers corresponding the automorphisms of ${\cal X}$ (Subsection 9.2). Finally, the inequality (2) enables us to find a $k$-local system of multipliers of ${\cal A}^{\star}$ for $k=\lfloor c\cdot a\rfloor$ for some constant $c=c(\varepsilon)$.
@@ 2.2. Coherent configurations
The number $|\Omega|$ is called the degree of ${\cal X}$. A coherent configuration ${\cal X}$ is said to be trivial if $S=S({\cal X})$ consists of $1_{\Omega}$ and its complement (unless $\Omega$ is a singleton), discrete if $S$ consists of singletons, and homogeneous or a scheme if $1_{\Omega}\in S$.
With any permutation group $G\leq\operatorname{Sym}(\Omega)$, one can associate a coherent configuration $\operatorname{Inv}(G)=(\Omega,S)$, where $S$ consists of the relations $(\alpha,\beta)^{G}=\{(\alpha^{g},\beta^{g}):\ g\in G\}$ for all $\alpha,\beta\in\Omega$. A scheme ${\cal X}$ is said to be regular if ${\cal X}=\operatorname{Inv}(G)$, where the group $G$ is regular, or equivalently, if $|\alpha s|=1$ for all $\alpha\in\Omega$ and $s\in S$.
@@ 2.3. Isomorphisms and algebraic isomorphisms
A combinatorial isomorphism or, briefly, isomorphism from a coherent configuration ${\cal X}=(\Omega,S)$ to a coherent configuration ${\cal X}^{\prime}=(\Omega^{\prime},S^{\prime})$ is defined to be a bijection $f:\Omega\rightarrow\Omega^{\prime}$ such that $S^{f}=S^{\prime}$. In this case, ${\cal X}$ and ${\cal X}^{\prime}$ are said to be isomorphic; the set of all isomorphisms from ${\cal X}$ to ${\cal X}^{\prime}$ is denoted by $\operatorname{Iso}({\cal X},{\cal X}^{\prime})$, and by $\operatorname{Iso}({\cal X})$ if ${\cal X}={\cal X}^{\prime}$. The group $\operatorname{Iso}({\cal X})$ contains a normal subgroup
$\operatorname{Aut}({\cal X})=\{f\in\operatorname{Sym}(\Omega):\ s^{f}=s\text{ for all }s\in S\}$
called the automorphism group of the coherent configuration ${\cal X}$.
A bijection $\varphi:S\rightarrow S^{\prime}$ is called an algebraic isomorphism from ${\cal X}$ to ${\cal X}^{\prime}$ if for all $r,s,t\in S$, we have
@@
In this case, $|\Omega^{\prime}|=|\Omega|$ and $|S|=|S^{\prime}|$. The set of all algebraic isomorphisms from ${\cal X}$ to ${\cal X}^{\prime}$ is denoted by $\operatorname{Iso}_{alg}({\cal X},{\cal X}^{\prime})$, and by $\operatorname{Aut}_{alg}({\cal X})$ if ${\cal X}={\cal X}^{\prime}$.
Every combinatorial isomorphism $f\in\operatorname{Iso}({\cal X},{\cal X}^{\prime})$ induces the algebraic isomorphism $\varphi_{f}:{\cal X}\to{\cal X}^{\prime},$ $s\mapsto s^{f}$; we put
$\operatorname{Iso}({\cal X},{\cal X}^{\prime},\varphi)=\{f\in\operatorname{Iso}({\cal X},{\cal X}^{\prime}):\ \varphi_{f}=\varphi\}$
and abbreviate $\operatorname{Iso}({\cal X},\varphi):=\operatorname{Iso}({\cal X},{\cal X},\varphi)$. Note that $\operatorname{Aut}({\cal X})=\operatorname{Iso}({\cal X},\operatorname{id}_{\cal X})$, where $\operatorname{id}_{\cal X}$ is the trivial (identical) algebraic automorphism of ${\cal X}$. If ${\cal X}$ is a regular scheme, then $\operatorname{Iso}({\cal X},{\cal X}^{\prime},\varphi)\neq\varnothing$ for any $\varphi$ (see [5, Theorem 2.3.33]).
@@ 2.4. Relations and fibers
Let ${\cal X}=(\Omega,S)$ be a coherent configuration. The elements of $S$ and of $S^{\cup}$ are called basis relations and relations of  ${\cal X}$. The basis relation containing a pair $(\alpha,\beta)\in{\bf 1}_{\Omega}$ is denoted by $r_{\cal X}(\alpha,\beta)$ or simply by $r(\alpha,\beta)$. The set of all relations of ${\cal X}$ is closed with respect to intersections, unions, and the dot product.
Any set $\Delta\subseteq\Omega$ such that $1_{\Delta}\in S$ is called a fiber of ${\cal X}$. In view of condition (CC1), the set $F=F({\cal X})$ of all fibers forms a partition of $\Omega$. Every basis relation is contained in the Cartesian product of two uniquely determined fibers.
Every algebraic isomorphism $\varphi\in\operatorname{Iso}_{alg}({\cal X},{\cal X}^{\prime})$ can be extended in a natural way to a poset isomorphism $S^{\cup}\to(S^{\prime})^{\cup}$, denoted also by $\varphi$; we have $1_{\Omega^{\prime}}=\varphi(1_{\Omega})$. This poset isomorphism respects the set theoretical operations and the dot product. Furthermore, $\varphi$ induces a poset isomorphism $F^{\cup}\to(F^{\prime})^{\cup}$ which preserves fibers. It takes a set $\Delta$ to the set $\Delta^{\varphi}$, and is defined by the equality $\varphi(1_{\Delta})=1_{\Delta^{\varphi}}$.
@@
An equivalence relation $e$ on a set $\Delta\subseteq\Omega$ is called a partial parabolic of the coherent configuration ${\cal X}$ if $e$ is a relation of ${\cal X}$; if, in addition, $\Delta=\Omega$, then $e$ is called a parabolic of ${\cal X}$. Note that the reflexive and transitive closure of any symmetric relation of ${\cal X}$ is a partial parabolic.
@@
respectively, where $s\in S$. Then the pairs ${\cal X}_{\Omega/e}=(\Omega/e,S_{\Omega/e})$ and ${\cal X}_{\Delta}=(\Delta,S_{\Delta})$ are coherent configurations called the quotient of ${\cal X}$ modulo $e$ and restriction of ${\cal X}$ to $\Delta$. Note that if $e=1_{\Delta}$ for some $\Delta\in F^{\cup}$, then ${\cal X}_{\Omega/e}={\cal X}_{\Delta}$.
Let $\varphi\in\operatorname{Iso}_{alg}({\cal X},{\cal X}^{\prime})$ be an algebraic isomorphism and $e$ a (partial) parabolic of ${\cal X}$. Then $\varphi(e)$ is a (partial) parabolic of ${\cal X}^{\prime}$ with the same number of classes. Moreover,
@@
(5) $\varphi_{\Omega/e}:{\cal X}_{\Omega/e}\to{\cal X}^{\prime}_{\Omega^{\prime}/e^{\prime}},\ s_{\Omega/e}\mapsto s^{\prime}_{\Omega^{\prime}/e^{\prime}},$
@@
(6) $\varphi_{\Delta,\Delta^{\prime}}:{\cal X}_{\Delta}\to{\cal X}^{\prime}_{\Delta^{\prime}},\ s_{\Delta}\mapsto s^{\prime}_{\Delta^{\prime}}.$
This algebraic isomorphism always exists for all $\Delta$ and $\Delta^{\prime}$ if ${\cal X}$ (and hence ${\cal X}^{\prime}$) is a scheme.
@@
Let ${\cal X}=(\Omega,S)$ and ${\cal X}^{\prime}=(\Omega^{\prime},S^{\prime})$ be two coherent configurations. Denote by $S\otimes S^{\prime}$ the set of all relations
@@
where $s\in S$ and $s^{\prime}\in S^{\prime}$. Then the pair ${\cal X}\otimes{\cal X}^{\prime}=(\Omega\times\Omega^{\prime},S\otimes S^{\prime})$ is a coherent configuration called the tensor product of ${\cal X}$ and ${\cal X}^{\prime}$. One can easily prove that $\operatorname{Aut}({\cal X}\times{\cal X}^{\prime})=\operatorname{Aut}({\cal X})\times\operatorname{Aut}({\cal X}^{\prime})$.
Assume that the coherent configurations ${\cal X}$ and ${\cal X}^{\prime}$ are schemes. Denote by $T$ and $T^{\prime}$ the sets of all relations
@@
respectively. Then the pair ${\cal X}\wr{\cal X}^{\prime}=(\Omega\times\Omega^{\prime},T\,\cup\,T^{\prime})$ is also a scheme. It is called the wreath product of ${\cal X}$ and ${\cal X}^{\prime}$. The mapping $\operatorname{Aut}({\cal X}\wr{\cal X}^{\prime})\to\operatorname{Aut}({\cal X}^{\prime})$ , $f\mapsto f^{\Omega^{\prime}},$ is an epimorphism the kernel of which is equal to the direct product of $|\Omega^{\prime}|$ copies of the group $\operatorname{Aut}({\cal X})$.

### v2

@@ 1. Introduction
For a positive integer $m$ and an arbitrary graph $\mathcal{G}$} (directed or undirected), the $m$-dimensional Weisfeiler-Leman algorithm (the $m$-dim $\operatorname{WL}$ algorithm) canonically constructs a special coloring of the $m$-tuples consisting of vertices of $\mathcal{G}$}. Two given graphs are declared isomorphic if the resulting colorings do not distinguish these graphs; in this case, the graphs are said to be $\operatorname{WL}_{m}$-equivalent. The algorithm works correctly if the $\operatorname{WL}_{m}$-equivalence of the input graphs implies that they are isomorphic. The $m$-dim $\operatorname{WL}$ algorithm is the naive canonical vertex classification in the sense of [2] if $m=1$, and is the classical Weisfeiler-Leman algorithm [30] if $m=2$. As one of the key ingredient, the $m$-dim $\operatorname{WL}$ algorithm was used in the Babai quasipolynomial algorithm testing isomorphism of graphs [3].
Following [15], the Weisfeiler-Leman dimension ($\operatorname{WL}$-dimension) $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{G}})$ of a graph $\mathcal{G}$} is defined to be the smallest positive integer $m$ such that any graph which is $\operatorname{WL}_{m}$-equivalent to the graph $\mathcal{G}$} is isomorphic to $\mathcal{G}$}. The isomorphism of any graph of order $n$ and $\operatorname{WL}$-dimension $m$ with any other graph can be tested in time $O(n^{m+1}\log n)$. Now the question we started with can be reformulated as follows (see [28]): whether there exists a positive integer $M$ such that $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{G}})\leq M$ for every circulant graph $\mathcal{G}$}.
It is known that the $\operatorname{WL}$-dimension of a general graph cannot be bounded from above by an absolute constant [4]. On the other hand, it was proved in [28] that if $\mathcal{G}$} is a circulant graph of prime power order, then $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{G}})\leq 3$. The main result of the present paper (Theorem 1.1 below) shows that the $\operatorname{WL}$-dimension of a general circulant graph can be arbitrary large. To formulate this result, we refer to Section 10 for the exact definition of $\varepsilon$-expanders, and given a positive integer $a$, denote by ${\mathfrak{B}}_{a}$ the set of all bipartite cubic $3$-connected graphs of order $a$.
@@
Following [32], we define the $\operatorname{WL}$-dimension of a circulant $\mathcal{G}$} with respect to the class of all circulant graphs as the smallest positive integer $m$ such that any circulant graph which is $\operatorname{WL}_{m}$-equivalent to $\mathcal{G}$} is isomorphic to $\mathcal{G}$}. It was proved in [32] that this “relative” $\operatorname{WL}$-dimension of any circulant of order $n$ is at most $O(\log n)$. The question of the tight asymptotic for the maximum relative $\operatorname{WL}$-dimension of circulants remains open. As Corollary 1.2 shows, it is between $O(\sqrt{\log n})$ and $O(\log n)$.
@@
We conclude the introduction with a brief discussion of the proof of Theorem 1.1. Our technique is based on theory of coherent configurations [5], see also Section 2. If $\mathcal{G}$} is a graph, then the output ${\mathcal{X}}=\operatorname{WL}({\mathcal{G}})$ of the $2$-dim $\operatorname{WL}$-algorithm applied to $\mathcal{G}$} can naturally be viewed as a coherent configuration. The concept of the $\operatorname{WL}$-dimension extends to coherent configurations such that
$\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{G}})=\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{X}}),$
@@
see Section 4. When the graph $\mathcal{G}$} is circulant, the coherent configuration $\mathcal{X}$} naturally corresponds to a Schur ring $\mathcal{A}$} over a cyclic group (a circulant S-ring), see Section 5. Now to find a circulant graph $\mathcal{G}$} with large $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{G}})$, we construct in Section 8 a circulant S-ring $\mathcal{A}$$^{\star}$ such that the corresponding coherent configuration $\mathcal{X}$$^{\star}$ has two properties: (a) $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{X}}^{\star})$ is large, and (b) ${\mathcal{X}}^{\star}=\operatorname{WL}({\mathcal{G}})$ for some circulant graph $\mathcal{G}$}. The construction of $\mathcal{A}$$^{\star}$ (see below) enables us to verify property (b) more or less easily, see Subsection 10.2.
@@
The construction of the S-ring $\mathcal{A}$$^{*}$ satisfying condition (a) is more subtle. First, we use a theory of multipliers in circulant S-rings, that was developed in [13, 14], see also Section 6. This theory is applied to express the $\operatorname{WL}_{m}$-equivalence of coherent configurations of circulant graphs in terms of $k$-local systems of multipliers of the corresponding circulant S-rings for $m<\sqrt{k}$, introduced and studied in Section 7. Thus our problem reduces to finding a circulant S-ring having a $k$-local system of multipliers for enough large $k$. An explicit construction of such an S-ring is given in Section 9 and is based on two ingredients defined in Section 8.
@@
The first ingredient is a class of coherent configurations, introduced and studied in [8], see Subsection 8.1. For any cubic graph $\mathcal{G}$}, one can define a coherent configuration ${\mathcal{X}}={\mathcal{X}}({\mathcal{G}})$ of this class so that: if $\mathcal{G}$} has $a$ vertices, then $\mathcal{X}$} has $4a$ points, and if $\mathcal{G}$} is a $3$-connected $\varepsilon$-expander, then
(2) $\dim_{\scriptscriptstyle\operatorname{WL}}({\mathcal{X}})\geq c^{\prime}\cdot a$
@@
The second ingredient is a class consisting of special (circulant) S-rings $\mathcal{A}$} over cyclic groups of orders $n$ satisfying the conditions of Theorem 1.1. The structure of $\mathcal{A}$} depends only on the number $n$, and is enough simple to describe all possible multipliers. In particular, if we fix a bipartite graph $\mathcal{G}$} from the previous paragraph, then isomorphisms of the coherent configuration $\mathcal{X}$} yield multipliers of $\mathcal{A}$}. Now the required S-ring $\mathcal{A}$$^{\star}$ is the algebraic fusion of the S-ring $\mathcal{A}$} with respects to the multipliers corresponding the automorphisms of $\mathcal{X}$} (Subsection 9.2). Finally, the inequality (2) enables us to find a $k$-local system of multipliers of $\mathcal{A}$$^{\star}$ for $k=\lfloor c\cdot a\rfloor$ for some constant $c=c(\varepsilon)$.
@@ 2.2. Coherent configurations
The number $|\Omega|$ is called the degree of $\mathcal{X}$}. A coherent configuration $\mathcal{X}$} is said to be trivial if $S=S({\mathcal{X}})$ consists of $1_{\Omega}$ and its complement (unless $\Omega$ is a singleton), discrete if $S$ consists of singletons, and homogeneous or a scheme if $1_{\Omega}\in S$.
With any permutation group $G\leq\operatorname{Sym}(\Omega)$, one can associate a coherent configuration $\operatorname{Inv}(G)=(\Omega,S)$, where $S$ consists of the relations $(\alpha,\beta)^{G}=\{(\alpha^{g},\beta^{g}):\ g\in G\}$ for all $\alpha,\beta\in\Omega$. A scheme $\mathcal{X}$} is said to be regular if ${\mathcal{X}}=\operatorname{Inv}(G)$, where the group $G$ is regular, or equivalently, if $|\alpha s|=1$ for all $\alpha\in\Omega$ and $s\in S$.
@@ 2.3. Isomorphisms and algebraic isomorphisms
A combinatorial isomorphism or, briefly, isomorphism from a coherent configuration ${\mathcal{X}}=(\Omega,S)$ to a coherent configuration ${\mathcal{X}}^{\prime}=(\Omega^{\prime},S^{\prime})$ is defined to be a bijection $f:\Omega\rightarrow\Omega^{\prime}$ such that $S^{f}=S^{\prime}$. In this case, $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ are said to be isomorphic; the set of all isomorphisms from $\mathcal{X}$} to $\mathcal{X}$$^{\prime}$ is denoted by $\operatorname{Iso}({\mathcal{X}},{\mathcal{X}}^{\prime})$, and by $\operatorname{Iso}({\mathcal{X}})$ if ${\mathcal{X}}={\mathcal{X}}^{\prime}$. The group $\operatorname{Iso}({\mathcal{X}})$ contains a normal subgroup
$\operatorname{Aut}({\mathcal{X}})=\{f\in\operatorname{Sym}(\Omega):\ s^{f}=s\text{ for all }s\in S\}$
called the automorphism group of the coherent configuration $\mathcal{X}$}.
A bijection $\varphi:S\rightarrow S^{\prime}$ is called an algebraic isomorphism from $\mathcal{X}$} to $\mathcal{X}$$^{\prime}$ if for all $r,s,t\in S$, we have
@@
In this case, $|\Omega^{\prime}|=|\Omega|$ and $|S|=|S^{\prime}|$. The set of all algebraic isomorphisms from $\mathcal{X}$} to $\mathcal{X}$$^{\prime}$ is denoted by $\operatorname{Iso}_{alg}({\mathcal{X}},{\mathcal{X}}^{\prime})$, and by $\operatorname{Aut}_{alg}({\mathcal{X}})$ if ${\mathcal{X}}={\mathcal{X}}^{\prime}$.
Every combinatorial isomorphism $f\in\operatorname{Iso}({\mathcal{X}},{\mathcal{X}}^{\prime})$ induces the algebraic isomorphism $\varphi_{f}:{\mathcal{X}}\to{\mathcal{X}}^{\prime},$ $s\mapsto s^{f}$; we put
$\operatorname{Iso}({\mathcal{X}},{\mathcal{X}}^{\prime},\varphi)=\{f\in\operatorname{Iso}({\mathcal{X}},{\mathcal{X}}^{\prime}):\ \varphi_{f}=\varphi\}$
and abbreviate $\operatorname{Iso}({\mathcal{X}},\varphi):=\operatorname{Iso}({\mathcal{X}},{\mathcal{X}},\varphi)$. Note that $\operatorname{Aut}({\mathcal{X}})=\operatorname{Iso}({\mathcal{X}},\operatorname{id}_{\mathcal{X}})$, where $\operatorname{id}_{\mathcal{X}}$ is the trivial (identical) algebraic automorphism of $\mathcal{X}$}. If $\mathcal{X}$} is a regular scheme, then $\operatorname{Iso}({\mathcal{X}},{\mathcal{X}}^{\prime},\varphi)\neq\varnothing$ for any $\varphi$ (see [5, Theorem 2.3.33]).
@@ 2.4. Relations and fibers
Let ${\mathcal{X}}=(\Omega,S)$ be a coherent configuration. The elements of $S$ and of $S^{\cup}$ are called basis relations and relations of  $\mathcal{X}$}. The basis relation containing a pair $(\alpha,\beta)\in{\bf 1}_{\Omega}$ is denoted by $r_{\mathcal{X}}(\alpha,\beta)$ or simply by $r(\alpha,\beta)$. The set of all relations of $\mathcal{X}$} is closed with respect to intersections, unions, and the dot product.
Any set $\Delta\subseteq\Omega$ such that $1_{\Delta}\in S$ is called a fiber of $\mathcal{X}$}. In view of condition (CC1), the set $F=F({\mathcal{X}})$ of all fibers forms a partition of $\Omega$. Every basis relation is contained in the Cartesian product of two uniquely determined fibers.
Every algebraic isomorphism $\varphi\in\operatorname{Iso}_{alg}({\mathcal{X}},{\mathcal{X}}^{\prime})$ can be extended in a natural way to a poset isomorphism $S^{\cup}\to(S^{\prime})^{\cup}$, denoted also by $\varphi$; we have $1_{\Omega^{\prime}}=\varphi(1_{\Omega})$. This poset isomorphism respects the set theoretical operations and the dot product. Furthermore, $\varphi$ induces a poset isomorphism $F^{\cup}\to(F^{\prime})^{\cup}$ which preserves fibers. It takes a set $\Delta$ to the set $\Delta^{\varphi}$, and is defined by the equality $\varphi(1_{\Delta})=1_{\Delta^{\varphi}}$.
@@
An equivalence relation $e$ on a set $\Delta\subseteq\Omega$ is called a partial parabolic of the coherent configuration $\mathcal{X}$} if $e$ is a relation of $\mathcal{X}$}; if, in addition, $\Delta=\Omega$, then $e$ is called a parabolic of $\mathcal{X}$}. Note that the reflexive and transitive closure of any symmetric relation of $\mathcal{X}$} is a partial parabolic.
@@
respectively, where $s\in S$. Then the pairs ${\mathcal{X}}_{\Omega/e}=(\Omega/e,S_{\Omega/e})$ and ${\mathcal{X}}_{\Delta}=(\Delta,S_{\Delta})$ are coherent configurations called the quotient of $\mathcal{X}$} modulo $e$ and restriction of $\mathcal{X}$} to $\Delta$. Note that if $e=1_{\Delta}$ for some $\Delta\in F^{\cup}$, then ${\mathcal{X}}_{\Omega/e}={\mathcal{X}}_{\Delta}$.
Let $\varphi\in\operatorname{Iso}_{alg}({\mathcal{X}},{\mathcal{X}}^{\prime})$ be an algebraic isomorphism and $e$ a (partial) parabolic of $\mathcal{X}$}. Then $\varphi(e)$ is a (partial) parabolic of $\mathcal{X}$$^{\prime}$ with the same number of classes. Moreover,
@@
(5) $\varphi_{\Omega/e}:{\mathcal{X}}_{\Omega/e}\to{\mathcal{X}}^{\prime}_{\Omega^{\prime}/e^{\prime}},\ s_{\Omega/e}\mapsto s^{\prime}_{\Omega^{\prime}/e^{\prime}},$
@@
(6) $\varphi_{\Delta,\Delta^{\prime}}:{\mathcal{X}}_{\Delta}\to{\mathcal{X}}^{\prime}_{\Delta^{\prime}},\ s_{\Delta}\mapsto s^{\prime}_{\Delta^{\prime}}.$
This algebraic isomorphism always exists for all $\Delta$ and $\Delta^{\prime}$ if $\mathcal{X}$} (and hence $\mathcal{X}$$^{\prime}$) is a scheme.
@@
Let ${\mathcal{X}}=(\Omega,S)$ and ${\mathcal{X}}^{\prime}=(\Omega^{\prime},S^{\prime})$ be two coherent configurations. Denote by $S\otimes S^{\prime}$ the set of all relations
@@
where $s\in S$ and $s^{\prime}\in S^{\prime}$. Then the pair ${\mathcal{X}}\otimes{\mathcal{X}}^{\prime}=(\Omega\times\Omega^{\prime},S\otimes S^{\prime})$ is a coherent configuration called the tensor product of $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$. One can easily prove that $\operatorname{Aut}({\mathcal{X}}\times{\mathcal{X}}^{\prime})=\operatorname{Aut}({\mathcal{X}})\times\operatorname{Aut}({\mathcal{X}}^{\prime})$.
Assume that the coherent configurations $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$ are schemes. Denote by $T$ and $T^{\prime}$ the sets of all relations
@@
respectively. Then the pair ${\mathcal{X}}\wr{\mathcal{X}}^{\prime}=(\Omega\times\Omega^{\prime},T\,\cup\,T^{\prime})$ is also a scheme. It is called the wreath product of $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$. The mapping $\operatorname{Aut}({\mathcal{X}}\wr{\mathcal{X}}^{\prime})\to\operatorname{Aut}({\mathcal{X}}^{\prime})$ , $f\mapsto f^{\Omega^{\prime}},$ is an epimorphism the kernel of which is equal to the direct product of $|\Omega^{\prime}|$ copies of the group $\operatorname{Aut}({\mathcal{X}})$.